OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=4, (7,1) The equation of the line in slope-intercept form is y = ?

OpenStudy (anonymous):

use the formula y-y1=m(x-x1) at point (7,1), where m=4 y-y1=m(x-x1) y-1=4(x-7) y=4(x-7)+1 y=4x-28+1 y=4x-27 ans